OpenStudy (anonymous):

A cone has a radius of 8 inches and a height of 12 inches. What is the exact volume of the cone? A. 32π in3 B. 96π in3 C. 256π in3 D. 768π in3

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

I think C

OpenStudy (anonymous):

nvm

Directrix (directrix):

I am here. Do you want to do this problem? @TeenWolfGirl

OpenStudy (anonymous):

i want to know if C is correct

Directrix (directrix):

Vcone = (1/3) * π * r² * h Vcone = (1/3) * π * 8² * 12 @TeenWolfGirl

OpenStudy (anonymous):

so yes?

Directrix (directrix):

That is what I and the Wolf got. @TeenWolfGirl

Directrix (directrix):

And, you.

Directrix (directrix):

Yes.